# A Generalization of von Neumann’s Reduction from the Assignment Problem to Zero-Sum Games

### v2

@@ 3 Reducing Linear Programs to Zero-Sum Games
  2. 2.
Incredibly, Dantzig’s reduction was nonetheless accepted and ascribed as complete for fifty years. Only then, a series of papers have worked towards a complete reduction. First, Adler (2013) modified Dantzig’s reduction and proposed a many-one reduction for the case that $A$, $b$, and $c$ are rational numbers. In the sequel, both Brooks and Reny (2023) and von Stengel (2023) noted that the output of Adler’s reduction was either a pair $(x^{*},y^{*})$ satisfying (1) in Proposition 2.1 or an assurance that there exists either $z^{*}$ or $w^{*}$ (possibly both) satisfying (2) or (3) in Proposition 2.1. Both papers introduce modifications adding a specific certificate of infeasibility that allows for solving $(P)$ as stated in Proposition 2.1.
The key idea of these papers is to modify the reduced game by Dantzig: Adler (2013) and von Stengel (2023) add an additional row while Brooks and Reny (2023) transform the game with elementary column operations. These reductions have in common the fact that the constraint matrix or vectors in the reduced games need large numbers, exponential in $m$ or nn. [^4] As an example, in the reduction by Adler (2013), the additional entry of the constraint vector is estimated as a number that is at least $m(\sqrt{m}\theta)^{m}+1$ where $\theta$ is the absolute value of the largest number contained in $A$ and $b$. A second way to solve linear programs explored by Adler (2013) is to use the original reduction by Dantzig and sequentially eliminate variables until the solution is in the right form. While this avoids large numbers, and is applicable to real numbers, it requires solving more than one zero-sum game. In other words, it is a Turing reduction rather than a many-one reduction. [^5] We refer to Section 7 for an in-depth discussion of the computational properties of our reduction.
4 Extending Von Neumann’s Reduction
In this section, we provide our extension of von Neumann’s reduction. First, we consider the case that the vectors $b>\mathbf{0}$ and $c>\mathbf{0}$, i.e., both are strictly positive. Next we will show how this enables us to tackle the case that the constraint matrix is nonnegative, i.e., $A\geq\mathbf{0}$.
4.1 Positive Objective and Constraint Vectors
The key technical step of our reduction is a rescaling of the linear programs $(P)$ and $(D)$; this is where we require $b>\mathbf{0}$ and $c>\mathbf{0}$, i.e., both vectors need to be strictly positive. We consider the diagonal matrices $B\in\mathbb{R}^{m\times m}$ and $C\in\mathbb{R}^{n\times n}$ defined by
[…]
We define the matrix $M=BAC\in\mathbb{R}^{m\times n}$ (its name already suggests that this will be the payoff matrix of our reduced zero-sum game) and a related pair of linear programs as follows.
@@ 4.1 Positive Objective and Constraint Vectors
The relationship with the original linear programs is captured as follows.
Proposition 4.1.
Assume that $b>\mathbf{0}$ and $c>\mathbf{0}$. The following statements are true:
  1. 1.
    1. (a)
      $\xi$ is feasible for $(P^{\prime})$ if and only if $C\xi$ is feasible for $(P)$,
    2. (b)
      $\eta$ is feasible for $(D^{\prime})$ if and only if $B\eta$ is feasible for $(D)$.
  2. 2.
    $(\xi^{*},\eta^{*})$ is an optimal primal-dual pair for $(P^{\prime})$ and $(D^{\prime})$ if and only if $(C\xi^{*},B\eta^{*})$ is an optimal primal-dual pair for $(P)$ and $(D)$.
  3. 3.
@@
Proof.
We prove the statements one by one.
  1. 1.
    Clearly, $\xi\geq 0$ if and only if $C\xi\geq 0$. Moreover, $M\xi\leq\mathbf{1}$ if and only if $A(C\xi)=B^{-1}M\xi\leq B^{-1}\mathbf{1}=b$. This implies (a). The proof of (b) is analogous.
  2. 2.
    From ($a$) and ($b$) we have that $\xi,\eta$ are feasible for $(P^{\prime}),(D^{\prime})$ if and only if $C\xi,B\eta$ are feasible for $(P),(D)$, respectively. In addition, $c^{\intercal}(C\xi)=\mathbf{1}^{\intercal}\xi$ and $b^{\intercal}(B\eta)=\mathbf{1}^{\intercal}\eta$. Applying proposition 2.1, completes the proof.
  3. 3.
    This follows from (1) and the observation that, because $c>\mathbf{0}$, $\mathbf{1}^{\intercal}\xi>0$ if and only if $\mathbf{1}^{\intercal}C\xi>0$. Hence, $\bar{\xi}\in\{\xi\;|\;M\xi\leq\mathbf{0},\;\xi\geq\mathbf{0},\;\mathbf{1}^{\intercal}\xi>0\}$ if and only if $C\bar{\xi}\in\{x\;|\;Ax\leq\mathbf{0},\;x\geq\mathbf{0},\;\mathbf{1}^{\intercal}x>0\}$.∎
Remark 4.2.
Proposition 4.1 covers cases (1) and (2) of Proposition 2.1, i.e., there exists an optimal primal-dual pair of solutions for $(P)$ or that $(P)$ is unbounded. Case (LABEL:it:inf), when $(P)$ and $(P^{\prime})$ are infeasible, never occurs under the assumption that $b>\mathbf{0}$, since in this case, $x=\mathbf{0}$ is a primal feasible solution.
Let $(G)$ be the zero-sum game with payoff matrix $M$ where the row-player wants to maximize his payoff and the column player wants to minimize her payoff. Recall that the maximin and minimax values of this game correspond to the optimal values of the linear programs introduced in Section 2:
$($ $\mbox{{Row}})$ $~($ $\mbox{{Col}})$
max $u$ min $v$
s.t. $M^{\intercal}p\geq\mathbf{1}u$ s.t. $Mq\leq\mathbf{1}v$
$\mathbf{1}^{\intercal}p=1,\;p\geq 0$ $\mathbf{1}^{\intercal}q=1,\;q\geq 0$
We have the following theorem.
@@
      If $v^{*}>0$, then $\left(\frac{1}{v^{*}}q^{*},\frac{1}{v^{*}}p^{*}\right)$ is an optimal primal-dual pair for $(P^{\prime})$ and $(D^{\prime})$.
@@
      If $w^{*}$ is an unboundedness certificate of $(P^{\prime})$, then $v^{*}\leq 0$.
@@
  2. 1-(b)
    Suppose that $v^{*}\leq 0$. Then, $Mq^{*}\leq\mathbf{1}v^{*}\leq\mathbf{0}$, $\mathbf{1}^{\intercal}q^{*}=1>0$, and $q^{*}\geq\mathbf{0}$. Hence, $q^{*}$ is an unboundedness certificate for $(P^{\prime})$.
  3. 2-(a)
    First, note that the optimal objective value of $(P^{\prime})$ is always strictly larger than 0, and, therefore, $\mathbf{1}^{\intercal}\xi^{*}>0$. Define $\hat{v}:=\frac{1}{\mathbf{1}^{\intercal}\xi^{*}}>0$. We have that $\mathbf{1}^{\intercal}(\hat{v}\eta^{*})=\mathbf{1}^{\intercal}(\hat{v}\xi^{*})=1$, $\hat{v}\eta^{*}\geq 0$, and $\hat{v}\xi^{*}\geq 0$, and, therefore, $\hat{v}\eta^{*}$ and $\hat{v}\xi^{*}$ are valid strategies.
@@
    Indeed, since $w^{*}\geq 0$, it holds that $\alpha w^{*}\geq 0$. In addition, $\mathbf{1}^{\intercal}q=\alpha\mathbf{1}^{\intercal}w^{*}=1$. Moreover, $Mq=\alpha Mw^{*}\leq\alpha\mathbf{0}=\mathbf{1}v$.
@@
In particular, the games by Dantzig and Adler are symmetric and have value zero, i.e., the value of the game has no information. [^6] While we regard the nonsymmetry of our reduced games as beneficial for interpreting the game, we acknowledge that reductions producing symmetric games have other, computational advantages. For instance, the symmetrization techniques by Gale et al. (1950) imply that algorithms for solving symmetric games can be applied to nonsymmetric games without significant further computational effort. The value of the reduced games by von Stengel as well as Brooks and Reny are zero if and only if there exists an optimal solution to the original LP. In contrast, the value of our reduced game has valuable information: it allows us to decide whether the LP admits an optimal solution and, if this is the case, it is the reciprocal of its optimal value. We believe this is a potent fact which may find a suitable application.
We conclude this subsection with a remark about how to interpret the value of the reduced game as a degree of (in)feasibility.
Remark 4.4.
@@
The primal-dual pair $(\mbox{{Row}},\mbox{{Col}})$ always has primal and dual feasible solutions, whereas $(D^{\prime})$ might be infeasible. Recall from Remark 2.2 that an infeasibility certificate for $(D^{\prime})$ is equivalent to an unboundedness certificate for $(P^{\prime})$, i.e., we ask whether there exists $w^{*}$ with $Mw^{*}\leq\mathbf{0}$, $w^{*}\leq 0$, and $\mathbf{1}^{\intercal}w^{*}>0$. By scaling $w^{*}$, the last condition can equivalently be written as $\mathbf{1}^{\intercal}w^{*}=1$. The LP $(\mbox{{Col}})$ only differs from these constraints by adding the minimum necessary slack to the constraints governed by $v$. Hence, the value of the reduced zero-sum game can be interpreted as a degree of feasibility, i.e., the minimum slack needed to guarantee feasibility of $(D^{\prime})$. Games with a nonpositive value indicate infeasibility of $(D^{\prime})$, and their exact value can then be seen as measuring a degree of infeasibility.
@@ 4.2 Nonnegative Constraint Matrix
Finally, we show that the correspondence described in Theorem 4.3 is applicable in a different setting as well. We demonstrate how to use the developed theory to deal with linear program pairs $(P,D)$ when $A\geq\mathbf{0}$ but $b$ and $c$ are not constrained to be positive.
The intuitive idea is to preprocess the linear program to satisfy the conditions that $b$ and $c$ are positive: If $b_{i}<0$, then the nonnegativity of $A$ immediately implies infeasibility. Otherwise, we want to delete all columns and rows that correspond to nonpositive entries of $b$ and $c$. If $b_{i}=0$, then $x_{j}=0$ whenever there exists $j\in[n]$ with $a_{ij}>0$. Hence, we can delete the $i$th row together with all such columns, and later on insert $x_{j}=0$ when reconstructing solutions. Similarly, $c_{j}\leq 0$ requires $x_{j}=0$ and we can omit such columns to transition to a smaller LP. We thus reach an LP of the original form, for which we can apply the correspondence of Section 4.1.
@@
We collect the key insights for the correctness of Algorithm 1 in the following theorem.
@@
    Suppose that $b\geq\mathbf{0}$.
@@
2-(b)
Consider the pair $(x^{*},y^{*})$ returned in algorithm 1. Clearly $x^{*}=\mathbf{0}$ is primal feasible. Moreover, $y^{*}\geq 0$. Now let $j\in[n]$. If $j\in J_{-}$, the $j$th constraint is satisfied. Otherwise, it holds that $j\in[n]\setminus J_{-}$ and we have already excluded the case where $a_{ij}=0$ for all $i\in[m]$. Hence, there exists $i\in[m]$ with $a_{ij}>0$. By definition of $\lambda$, it holds that $\lambda\geq\frac{c_{j}}{a_{ij}}$. Therefore, $(A^{\intercal}y^{*})_{j}\geq a_{ij}y^{*}_{i}=a_{ij}\lambda\geq a_{ij}\frac{c_{j}}{a_{ij}}=c_{j}$. Hence, $y^{*}$ is dual feasible. Finally, it holds that $b^{\intercal}y^{*}=\mathbf{0}=c^{\intercal}x^{*}$ where the first inequality holds by definition of $y^{*}$ (whose components are only nonzero for indices in $I_{0}$). Together, $(x^{*},y^{*})$ is an optimal primal-dual pair.
3.
@@
Moreover, since $\hat{b}\geq 0$, we can apply part 2-(a) to conclude that (P^)($\hat${P}) is bounded if $\hat{A}$ does not contain 0-columns (which is a stronger condition than only requiring certain nonzero columns). Let $j\in[n]\setminus(J_{0}\cup J_{-})$. We claim that there exists $i\in[m]\setminus I_{0}$ with $a_{ij}>0$. Assume that this was not the case, i.e., $a_{ij}=0$ for all $i\in[m]\setminus I_{0}$. Note that since $j\notin J_{0}$, we also have that $a_{ij}=0$ for all $i\in I_{0}$. However, this would imply that we have identified a $j\in[n]\setminus J_{-}$ such that $a_{ij}=0$ for all $i\in[m]$, contradicting that Algorithm 1 has reached algorithm 1. Hence, (P^)($\hat${P}) is feasible and bounded, and, therefore, (D^)($\hat${D}) is feasible as well. By strong LP duality, we have the existence of an optimal primal-dual pair.
@@
5 Application and Interpretation
In this section, we want to demonstrate the applicability of our reductions. We start by deepening our understanding of von Neumann’s hide-and-seek game in the context of our more general reduction. We then discuss playing the reduced game for any LP with positive $b$ and $c$. Finally, we provide an interpretation of the reduced game in relationship to its original LP, where we interpret the players’ strategies as “playing” the primal and dual LP.
5.1 Von Neumann’s Correspondence for the Assignment Problem
Originally, von Neumann had designed his reduction for the special case where the source LP corresponds to the (fractional) assignment problem. The assignment problem consists of $n$ workers and $n$ jobs and it is assumed that assigning worker $i$ to job $j$ creates a positive surplus $\mu_{ij}>0$. A fractional assignment of workers to jobs is a nonnegative matrix $(x_{ij})$, where
  • for all $1\leq i\leq n$, it holds that $\sum_{j=1}^{n}x_{ij}\leq 1$, and
  • for all $1\leq j\leq n$, it holds that $\sum_{i=1}^{n}x_{ij}\leq 1$.
The constraints ensure that each worker is assigned no more than the fractions equivalent to one job, and no job is assigned more than one worker. The goal is to find a fractional assignment that maximizes the expected surplus. [^7] Due to the famous algorithm by Birkhoff (1946), an optimal fractional assignment can be decomposed into perfect matchings of the same value. Hence, finding an optimal fractional assignment is sufficient to find an optimal integral assignment. In fact, the paper by von Neumann (1953) is more famous as a reference for the Birkhoff-von Neumann Decomposition Theorem of bistochastic matrices rather than for introducing the hide-and-seek game. This is equivalent to solving the linear program $(\mbox{{Asg}})$ whose dual is given by $(\overline{\mbox{{Asg}}})$.
$(\mbox{{Asg}})$
$\text{max}\qquad\sum_{i,j=1}^{n}\mu_{ij}x_{ij}$
$\text{s.t. }\qquad\sum_{j=1}^{n}x_{ij}$ $\leq 1\ \forall\,i\in[n]$
$\qquad\sum_{i=1}^{n}x_{ij}$ $\leq 1\ \forall\,j\in[n]$
$x_{ij}$ $\geq 0\ \forall\,i,j\in[n]$
$(\overline{\mbox{{Asg}}})$
$\text{min}\qquad\sum_{i=1}^{n}u_{i}+\sum_{j=1}^{n}v_{j}$
$\text{s.t. }\qquad u_{i}+v_{j}$ $\geq\mu_{ij}\ \forall\,i,j\in[n]$
$u_{i}$ $\geq 0\ \forall\,i\in[n]$
$v_{j}$ $\geq 0\ \forall\,j\in[n]$
(Asg) is a linear program of the form $(P)$ where the objective vector is strictly positive (containing the entries $\mu_{ij}$), the constraint matrix contains only zeros and ones, and the right hand side is the all-ones vector. Its dual contains variables $u_{i}$ and $v_{j}$ corresponding to workers and jobs of the assignment problem.
Now, matrix $B$ for $(\mbox{{Asg}})$ is the identity matrix, and matrix $C$ is the diagonal matrix of dimension $n^{2}\times n^{2}$ with entries $\frac{1}{\mu_{ij}}$. Rescaling of the linear program analogous for obtaining $(P^{\prime},D^{\prime})$ then yields the following linear programs.
$(\mbox{{Asg}}^{\prime})$
$\text{max}\qquad\sum_{i,j=1}^{n}x_{ij}$
$\text{s.t. }\qquad\sum_{j=1}^{n}\frac{1}{\mu_{ij}}x_{ij}$ $\leq 1\ \forall\,i\in[n]$
$\qquad\sum_{i=1}^{n}\frac{1}{\mu_{ij}}x_{ij}$ $\leq 1\ \forall\,j\in[n]$
$x_{ij}$ $\geq 0\ \forall\,i,j\in[n]$
$(\overline{\mbox{{Asg}}}^{\prime})$
$\text{min}\qquad\sum_{i=1}^{n}u_{i}+\sum_{j=1}^{n}v_{j}$
$\text{s.t. }\qquad\frac{1}{\mu_{ij}}u_{i}+\frac{1}{\mu_{ij}}v_{j}$ $\geq 1\ \forall\,i,j\in[n]$
$u_{i}$ $\geq 0\ \forall\,i\in[n]$
$v_{j}$ $\geq 0\ \forall\,j\in[n]$
The constraint matrix of $(\mbox{{Asg}}^{\prime})$ can now be used to define a zero-sum game, the hide-and-seek game by von Neumann. The game matrix has dimension $2n\times n^{2}$, with only $2n^{2}$ of its $2n^{3}$ entries being nonzero. The nonzero entries correspond exactly to the surplus matrix of the assignment problem.
Recall that we can view the hide-and-seek game as the row and column players picking a vertex and edge of the assignment graph, respectively. The column player tries to hide in the edge and the row player tries to search edges by inspecting end points. Our correspondence explains why the payoffs of the hide-and-seek game are the inverse of the matching surplus of the respective cell. This is due to the rescaling of the linear program to rewrite it in the right form for games. Moreover, we can also revisit von Neumann’s correspondence of maximum weight matchings and strategies of the column player. His main theorem says that the maximin strategies of the column player (von Neumann calls them optimal strategies) are obtained by permutation matrices corresponding to maximum weight matchings. More precisely, if we view a permutation as a bijective function $\sigma\colon[n]\to[n]$, then the maximin strategies are convex combinations of strategies with support determined by permutations $\sigma$ that maximize $\sum_{i=1}^{n}\mu_{i\sigma(i)}$. Within the support of the matching specified by a permutation, the probabilities $p$ must be distributed such that $p_{1}\frac{1}{\mu_{1\sigma(1)}}=\dots=p_{n}\frac{1}{\mu_{n\sigma(n)}}$. [^8] This proportion is captured by von Neumann in the analysis of a so-called $1$-dimensional game that corresponds to the analysis of game play given a fixed permutation matrix (or, equivalently, perfect matching). Hence, it is again the rescaling to obtain $(P^{\prime},D^{\prime})$ that tells how feasible solutions of the primal will be skewed to obtain maximin strategies of the game. Once we consider linear programs of the form $(P^{\prime},D^{\prime})$, the correspondence of optimal primal-dual pairs and maximin strategies requires no more skewing but simply multiplication with a scalar. Put differently, the scaled matrix leads to a uniform scale for the payoffs.
5.2 Production Planning Game

### v3

@@ 3.1 Positive Objective and Constraint Vectors
We define the matrix $M=BAC\in\mathbb{R}^{m\times n}$ (its name already suggests that this will be the payoff matrix of our reduced zero-sum game) and a related pair of LPs as follows.
@@
It is easy to see that this change of variables by scaling merely leads to a rescaling in the solutions of the original pair of LPs:
  1. (i)
    $\xi$ (resp., $\eta$) is feasible for $(P^{\prime})$ (resp., $(D^{\prime})$) if and only if $C\xi$ (resp., $B\eta$) is feasible for $(P)$ (resp., $(D)$)
  2. (ii)
    $(\xi^{*},\eta^{*})$ is an optimal primal-dual pair for $(P^{\prime},D^{\prime})$ if and only if $(C\xi^{*},B\eta^{*})$ is an optimal primal-dual pair for $(P,D)$.
  3. (iii)
@@
We remark that $(P)$ and $(P^{\prime})$ are never infeasible when $b>\mathbf{0}$, since in this case, $x=\mathbf{0}$ is a feasible solution.
Now, let $(G)$ be the zero-sum game with payoff matrix $M$ where the row-player wants to maximize his payoff and the column player wants to minimize her payoff. Recall that the maximin and minimax values of this game correspond to the optimal values of the LPs $(\mbox{{Row}})$ and $(\mbox{{Col}})$ defined in Section 2. We have the following theorem. Note that while it establishes a correspondence between zero-sum games and LP pairs of the type $(P^{\prime},D^{\prime})$, it applies to LP pairs of the type $(P,D)$ whenever $b>\mathbf{0}$ and $c>\mathbf{0}$, i.e., the rescaling can be performed.
@@
      If $v^{*}>0$, then $\left(\frac{1}{v^{*}}q^{*},\frac{1}{v^{*}}p^{*}\right)$ is an optimal primal-dual pair for $(P^{\prime},D^{\prime})$.
@@
      If $(P^{\prime})$ is unbounded, then $v^{*}\leq 0$.
@@
  2. 1-(b).
    Suppose that $v^{*}\leq 0$. Then, $Mq^{*}\leq v^{*}\mathbf{1}\leq\mathbf{0}$, $\mathbf{1}^{\intercal}q^{*}=1>0$, and $q^{*}\geq\mathbf{0}$. Hence, $q^{*}$ is an unboundedness certificate for $(P^{\prime})$.
  3. 2-(a).
    First, note that the optimal objective value of $(P^{\prime})$ is always strictly larger than 0, and, therefore, $\mathbf{1}^{\intercal}\xi^{*}>0$. Define $\hat{v}:=\frac{1}{\mathbf{1}^{\intercal}\xi^{*}}>0$. We have that $\mathbf{1}^{\intercal}(\hat{v}\eta^{*})=\mathbf{1}^{\intercal}(\hat{v}\xi^{*})=1$, $\hat{v}\eta^{*}\geq\mathbf{0}$, and $\hat{v}\xi^{*}\geq\mathbf{0}$, and, therefore, $\hat{v}\eta^{*}$ and $\hat{v}\xi^{*}$ are valid strategies.
@@
    Indeed, since $w^{*}\geq\mathbf{0}$, it holds that $\alpha w^{*}\geq\mathbf{0}$. In addition, $\mathbf{1}^{\intercal}q=\alpha\mathbf{1}^{\intercal}w^{*}=1$. Moreover, $Mq=\alpha Mw^{*}\leq\alpha\mathbf{0}=v\mathbf{1}$.
@@
We conclude this subsection with remarks about the interpretation of our reduction.
Linear Programs in Game Form
@@
Once the rescaling is applied, we obtain a correspondence by directly using the constraint matrix as the game matrix of a zero-sum game. In other words, LPs of the form $(P^{\prime},D^{\prime})$ can be interpreted as linear programs in game form. Hence, von Neumann’s hide-and-seek game results from the realization that the assignment problem can be written in game form. And more generally, we observe that any linear program pair of the form $(P,D)$ with $b>\mathbf{0}$ and $c>\mathbf{0}$ can be transformed to game form, as well.
Unlike in previous reductions, the column and row players in our reduction have different roles corresponding to the primal and dual linear programs, respectively. Each fixed strategy (including a nonoptimal one) gives a guarantee to a player, i.e., a minimum reward for the row player against adversarial play of the column player or a maximum payment by the column player against an adversarial row player. These guarantees are obtained by solving $(\mbox{{Row}})$ and $(\mbox{{Col}})$ after fixing strategies $p$ and $q$, respectively. When players play optimal strategies, they get their optimal guarantee, which is exactly the maximin or minimax value.
In fact, one can view playing the game of an LP in game form as playing the linear program. Any strategy guaranteeing the column player nonpositive payoff (not only the maximin strategy) gives rise to an unboundedness certificate as in Part 1-(b) of the proof of Theorem 3.1. Moreover, strategies guaranteeing the row player a positive payoff correspond to dual feasible points. Once again, this works even for nonoptimal strategies, but can then lead to feasible points with a suboptimal objective value. Hence, players compete by proposing (good) feasible points of their respective LPs, whenever this is possible. This can be viewed as playing an LP. The reduced zero-sum game yields a reasonable outcome of playing, even for suboptimal or infeasible strategies. In the special case, where both players play maximin strategies, we obtain an optimal primal-dual pair.
Value of Game as Degree of Infeasibility
@@
The games by Dantzig and Adler are symmetric and have value zero, i.e., the value of the game has no information. [^3] While we regard the nonsymmetry of our reduced games as beneficial for interpreting the game, we acknowledge that reductions producing symmetric games have other, computational advantages. For instance, the symmetrization techniques by Gale et al. (1950) imply that algorithms for solving symmetric games can be applied to nonsymmetric games without significant further computational effort. Moreover, the value of the reduced games by von Stengel as well as Brooks and Reny are zero if and only if there exists an optimal solution to the original LP. In contrast, the value of our reduced game has valuable information: first of all, it allows us to decide whether the LP admits an optimal solution and, if so, it is the reciprocal of its optimal value.
@@
Additionally, the primal-dual pair $(\mbox{{Row}},\mbox{{Col}})$ always has primal and dual feasible solutions, whereas $(D^{\prime})$ might be infeasible. Recall from Remark 2.2 that an infeasibility certificate for $(D^{\prime})$ is equivalent to an unboundedness certificate for $(P^{\prime})$, i.e., we ask whether there exists $w^{*}$ with $Mw^{*}\leq\mathbf{0}$, $w^{*}\geq\mathbf{0}$, and $\mathbf{1}^{\intercal}w^{*}>0$. By scaling $w^{*}$, the last condition can equivalently be written as $\mathbf{1}^{\intercal}w^{*}=1$. The LP $(\mbox{{Col}})$ only differs from these constraints by adding the minimum necessary slack to the constraints governed by $v$. Hence, the value of the reduced zero-sum game can be interpreted as a degree of feasibility, i.e., the minimum slack needed to guarantee feasibility of $(D^{\prime})$. Games with a nonpositive value indicate infeasibility of $(D^{\prime})$, and their exact value can then be seen as measuring a degree of infeasibility.
@@ 3.2 Nonnegative Constraint Matrix
Finally, we show that the developed theory in Section 3.1 also allows us to deal with linear program pairs $(P,D)$ when $A\geq\mathbf{0}$ but $b$ and $c$ are not constrained to be positive.
The intuitive idea is to preprocess the LP to satisfy the conditions that $b$ and $c$ are positive: If $b_{i}<0$, then the nonnegativity of $A$ immediately implies infeasibility. Otherwise, we want to delete all columns and rows that correspond to nonpositive entries of $b$ and $c$. If $b_{i}=0$, then $x_{j}=0$ whenever there exists $j\in[n]$ with $a_{ij}>0$. Hence, we can delete the $i$th row together with all such columns, and later on insert $x_{j}=0$ when reconstructing solutions. Similarly, $c_{j}\leq 0$ requires $x_{j}=0$ and we can omit such columns to transition to a smaller LP. We thus reach a reduced LP with $b>\mathbf{0}$ and $c>\mathbf{0}$, for which we can apply the correspondence of Section 3.1.
@@
  2. 3.
    Otherwise $(P,D)$ admits an optimal primal-dual solution. There are two cases for Algorithm 1:
@@
  3. 3-(a).
    Consider the pair $(x^{*},y^{*})$ returned in algorithm 1. Clearly $x^{*}=\mathbf{0}$ is primal feasible. Moreover, $y^{*}\geq\mathbf{0}$. Now let $j\in[n]$. If $j\in J_{-}$, the $j$th constraint is satisfied. Otherwise, it holds that $j\in[n]\setminus J_{-}$ and we have already excluded the case where $a_{ij}=0$ for all $i\in[m]$. Hence, there exists $i\in[m]$ with $a_{ij}>0$. By definition of $\lambda$, it holds that $\lambda\geq\frac{c_{j}}{a_{ij}}$. Therefore, $(A^{\intercal}y^{*})_{j}\geq a_{ij}y^{*}_{i}=a_{ij}\lambda\geq a_{ij}\frac{c_{j}}{a_{ij}}=c_{j}$. Hence, $y^{*}$ is dual feasible. Finally, it holds that $b^{\intercal}y^{*}=\mathbf{0}=c^{\intercal}x^{*}$ where the first inequality holds by definition of $y^{*}$ (whose components are only nonzero for indices in $I_{0}$). Together, $(x^{*},y^{*})$ is an optimal primal-dual pair.
  4. 3-(b).
@@
    Moreover, since $\hat{b}\geq\mathbf{0}$, we can apply part 2-(a) to conclude that (P^)($\hat${P}) is bounded if $\hat{A}$ does not contain 0-columns (which is a stronger condition than only requiring certain nonzero columns). Let $j\in[n]\setminus(J_{0}\cup J_{-})$. We claim that there exists $i\in[m]\setminus I_{0}$ with $a_{ij}>0$. Assume that this was not the case, i.e., $a_{ij}=0$ for all $i\in[m]\setminus I_{0}$. Note that since $j\notin J_{0}$, we also have that $a_{ij}=0$ for all $i\in I_{0}$. However, this would imply that we have identified a $j\in[n]\setminus J_{-}$ such that $a_{ij}=0$ for all $i\in[m]$, contradicting that Algorithm 1 has reached algorithm 1. Hence, (P^)($\hat${P}) is feasible and bounded, and, therefore, (D^)($\hat${D}) is feasible as well. By strong LP duality, we have the existence of an optimal primal-dual pair.
@@
4 Discussion
The question of generalizing von Neumann’s result (von Neumann, 1953) to an arbitrary primal-dual pair of LPs is clearly an important one. Our work explores an extension based on rescaling, the crucial idea that helps define the payoffs in the hide-and-seek game. While at first glance this seems only possible if $b>\mathbf{0}$ and $c>\mathbf{0}$, we managed to reduce the case $A\geq\mathbf{0}$ to this case as well. Going further may well require a new approach.
We conclude our paper with discussions tying up a few loose ends.
4.1 Production Planning Games
We want to demonstrate the applicability of our reductions by discussing an interpretation of the reduced game for any LP with positive $b$ and $c$ and demonstrate how von Neumann’s hide-and-seek game is a special case.